SIR PLZ ANSWER AS FAST AS POSSIBLE !!!!!!!!!!!!!!!!1 
I WILL BE HELPFULL
The sum of n terms of the following series; 13 + 33 + 53 + 73 + .... is
(A) n2(2n2 – 1)
 (B) n3(n – 1) 
(C) n3 + 8n + 4 
(D) 2n4 + 3n2

Grade:12th pass

4 Answers

Arun
6 years ago
Dear Sohan
 
Please uae hit and trial. It will save your time.
As put n = 1
Hence sum should be equal to 13. Which is option C
Samyak Jain
5 years ago
Let S denote the sum of n terms of the series Tbe the nth term of the series.
S = 13 + 33 + 53 + 73 + ....... + Tn                                      ..................(1)
S =          13 + 33 + 53 + ....... + Tn –1 + Tn                    ….............(2)
Subtract (2) from (1)
0 = 13 + (20 + 20 + 20 + …........ n – 1 times) –  Tn
\therefore  Tn = 13 + (n – 1) * 20  =  20n – 20 + 13  =  20n – 7
\therefore  S  =  \sum_{}^{}n=1n=n  Tn   =    ∑ 20n – ∑7 
         = [20n(n + 1) / 2]  –   7n
         = 10n(n+1) – 7n   
         = 10n+ 10n – 7n
         = 10n+ 3n
S  =  n (10n + 3) is the correct answer.
All the options given are incorrect  !!!!
Option C is valid only for n=1.
Govind
5 years ago
Please uae hit and trial. It will save your time.As put n = 1Hence sum should be equal to 13. Which is option C orLet S denote the sum of n terms of the series Tn be the nth term of the series.S = 13 + 33 + 53 + 73 + ....... + Tn ..................(1)S = 13 + 33 + 53 + ....... + Tn –1 + Tn ….............(2)Subtract (2) from (1)0 = 13 + (20 + 20 + 20 + …........ n – 1 times) – Tn\therefore Tn = 13 + (n – 1) * 20 = 20n – 20 + 13 = 20n – 7\therefore S = \sum_{}^{}n=1n=n Tn = ∑ 20n – ∑7 = [20n(n + 1) / 2] – 7n = 10n(n+1) – 7n = 10n2 + 10n – 7n = 10n2 + 3nS = n (10n + 3) is the correct answer.All the options given are incorrect !!!!Option C is valid only for n=1.
Om
5 years ago
Dear student please check the question once more
All the options are incorrect 
We can confirm this either  by hit and trial or by solving it
In c option only 13 is correct 
I hope this answer helps you